Find the z critical value(ta/2,n-1) to construct a 92% large sample z confidence interval.
Answer #1

solution:

At 92% confidence level the z is ,

= 1 - 92% = 1 - 0.92 = 0.08

/ 2 = 0.08 / 2 = 0.04

Z/2 = Z0.04 = 1.751 ( Using z table ( see the 0.04 value in standard normal (z) table corresponding z value is 1.751 )